# Evaluating the Defense Potential of Machine Unlearning against Membership Inference Attacks

### v3

@@ 6 Results
6.1 CIFAR-10 Dataset
6.1.1 NegGrad Algorithm
We begin by examining the NegGrad algorithm applied to the CIFAR-10 dataset, which yields several insights into the unlearning process. A key highlight here is that when the forget accuracy reaches zero, it signifies that the model has successfully forgotten the specified data, as illustrated in Fig 3. Additionally, during the unlearning process, we observe a decline in test accuracies. This highlights a trade-off inherent in the algorithm: while the unlearning process effectively achieves perfect forgetting, it comes at the cost of reduced overall model performance.
Figure 3: ML Model Accuracy on CIFAR-10 Dataset / Unlearning Algorithm NegGrad
Furthermore, the Fig 4 illustrates the accuracy scores of the MIA attack on the forget and retained data, as the model undertakes the unlearning process per epoch. The accuracy for the forget data (solid blue line) shows a consistent decline, indicating that the unlearning process is effectively reducing the model’s ability to recall the forget data points, which aligns with the intended goal of making the model forget specific data. Conversely, the accuracy for the retained data (dashed orange line) also decreases over time, though at a slower pace initially. This outcome was expected, as the unlearning algorithm not only reduces the model’s ability to recall forget data but also impacts overall model performance, which in turn lowers the effectiveness of MIA attacks on both forget and retained data.
Figure 4: MIA Accuracy on CIFAR-10 Dataset / Unlearning Algorithm NegGrad
6.1.2 SCRUB Algorithm
On the other hand, the scrub algorithm applied to the CIFAR-10 dataset did not demonstrate the same functionality as the NegGrad approach. As depicted in Fig 5, the forget accuracy remained stable, showing no significant decline across epochs. Additionally, in the context of MIAs (see Fig. 6), we observed a slight decrease in the forget accuracy score. This suggests that the unlearning process may have some impact on mitigating the effectiveness of MIA attacks.
Figure 5: ML Model Accuracy on CIFAR-10 Dataset / Unlearning Algorithm SCRUB
Figure 6: MIA Accuracy on CIFAR-10 Dataset / Unlearning Algorithm SCRUB
6.1.3 SFTC Algorithm
In Figure 7 the results show that the SFTC algorithm on the CIFAR-10 dataset maintains high performance on the training data, with stable test and validation accuracies. However, the forget accuracy remains high, suggesting that the algorithm is not effectively unlearning the forget data. Compared to the other two algorithms, SFTC performs better in terms of maintaining model utility and generalization ability, making it the best option among the three for balancing unlearning and performance.
Figure 7: ML Model Accuracy on CIFAR-10 Dataset / Unlearning Algorithm SFTC
Figure 8 shows the MIA accuracy trends for SFTC, where the forget accuracy steadily decreases while the retain accuracy remains relatively stable. This result suggests that SFTC effectively mitigates MIA risks by successfully unlearning forget data while maintaining model performance on retained data. Furthermore, this algorithm outperforms the others in mitigating MIA attacks, as evidenced by the lower MIA accuracy on the retained dataset compared to the other algorithms. Only the NegGrad algorithm achieved better (lower) MIA attack accuracies, but it does so at the cost of also degrading the performance on the retained dataset. This makes SFTC more effective overall, as it balances successful unlearning with maintaining model performance on the retained data.
Figure 8: MIA Model Accuracy on CIFAR-10 Dataset / Unlearning Algorithm SFTC
6.2 MuFac Dataset
6.2.1 NegGrad Algorithm
The NegGrad algorithm performs significantly better on the MuFac dataset compered to CIFAR-10, as demonstrated by the Figure 9. This can be attributed to the simpler and more structured nature of the MuFac dataset, which likely consists of facial images with less variability between classes. Furthermore, the near-identical accuracy trends in both the training and test sets during the epochs for the MuFac dataset suggest that the model maintained stability. Lastly, the clear gap between the training accuracy, which remains at nearly 1.0, and the test accuracy, which hovers around 0.6, indicates significant overfitting.
Figure 9: ML Model Accuracy on MuFac Dataset / Unlearning Algorithm NegGrad
Figure 10 presents the MIA accuracy trends, where the forget accuracies decline sharply over the epochs, indicating that NegGrad is effectively unlearning the forget data with respect to MIA attacks. This declining trend in forget accuracies aligns with the goal of reducing the model’s ability to recall forget data. Meanwhile, the retain accuracies remain relatively stable throughout the unlearning process, suggesting that the model’s ability to retain and accurately classify non-forget data is preserved. This balance shows that NegGrad is successfully mitigating MIA risks for forget data without significantly compromising the accuracy of retained data.
Figure 10: MIA Model Accuracy on MuFac Dataset / Unlearning Algorithm NegGrad
6.2.2 SCRUB Algorithm
Figure 11 shows the SCRUB algorithm’s performance on MuFac, where the forget accuracy shows a minor decline but remains relatively stable throughout the epochs. The validation and test accuracies also remain stable, indicating that the algorithm preserves model performance but may not be as effective in unlearning forget data.
Figure 11: ML Model Accuracy on MuFac Dataset / Unlearning Algorithm SCRUB
In Figure 12, the MIA accuracy trends for the SCRUB algorithm on the MuFac dataset are shown. The forget accuracies fluctuate considerably across the epochs but remain relatively high, hovering around 0.62-0.64. This suggests that the SCRUB algorithm is less effective at consistently unlearning the forget data, as the model still maintains a relatively high ability to recall forget information. Meanwhile, the retain accuracies also show fluctuations but remain slightly lower than the forget accuracies, mostly around 0.60-0.62. This indicates that the model’s ability to retain data is not significantly stronger than its ability to recall forget data, which suggests a trade-off between unlearning forget data and maintaining the retained data.
Figure 12: MIA Model Accuracy on MuFac Dataset / Unlearning Algorithm SCRUB
6.2.3 SFTC Algorithm
In Figure 13 the SFTC algorithm applied to the MuFac dataset shows high training accuracy, consistently close to 1.0, indicating that the model fits the training data extremely well. However, the test accuracy fluctuates around 0.3, revealing a significant gap between the training and test performance, which indicates substantial overfitting. Furthermore, the forget accuracy hovers around 0.85, suggesting that the algorithm has not been fully effective in making the model forget the specified data, as it retains a high level of accuracy on the forget data.
Figure 13: ML Model Accuracy on MuFac Dataset / Unlearning Algorithm SFTC
In Figure 14 we observe the MIA attack accuracies for the SFTC algorithm on the MuFac dataset as the model unlearns over time. The forget accuracies remain relatively stable, fluctuating between 0.60 and 0.65 across the epochs. This indicates that the model’s vulnerability to MIA attacks on forget data is reducing, which is aligned with the unlearning objective, but the algorithm is not fully effective in completely mitigating MIA attacks on forget data. In contrast, the retain accuracies are higher, fluctuating between 0.70 and 0.75, suggesting that the model remains more vulnerable to MIA attacks on retained data. This higher and more variable accuracy for retained data suggests that while the SFTC algorithm successfully diminishes the model’s ability to recall forget data (as seen by the lower MIA accuracy), it still struggles to maintain a consistent level of resistance to MIA attacks on retained data.
Figure 14: MIA Model Accuracy on MuFac Dataset / Unlearning Algorithm SFTC
6.3 PURCHASE Dataset
6.3.1 NegGrad Algorithm
In Figure 15, we observe the performance of the NegGrad algorithm on the Purchase dataset, showing trends that closely mirror those observed with the CIFAR-10 dataset. The training accuracy and forget accuracy both start high, around 0.8, but decline steadily over the epochs, reaching approximately 0.1 by the end. This sharp decline indicates that the algorithm is effectively unlearning the specified data, as the model gradually loses its ability to recall the forget information, similar to its behavior on CIFAR-10. Furthermore, the results on the Purchase dataset confirm the pattern observed with CIFAR-10: the NegGrad algorithm effectively forgets targeted data, but this unlearning process leads to a notable decline in test and validation accuracies, indicating compromised generalization. This highlights the inherent trade-off in the NegGrad algorithm, where successful unlearning is achieved at the expense of the model’s ability to retain its performance on retained and unseen data.
Figure 15: ML Model Accuracy on Purchase Dataset / Unlearning Algorithm NegGrad
In Figure 16 the MIA attack accuracies for the NegGrad algorithm on the Purchase dataset demonstrate how the model’s vulnerability to MIA attacks changes during the unlearning process. The forget accuracies steadily decrease from approximately 0.58 to 0.48 over the course of 50 epochs, indicating the algorithm’s success in making the model forget specific data points. This declining trend highlights NegGrad’s effectiveness in reducing the model’s ability to recall forget data, thereby achieving the intended unlearning objective. However, as seen with other datasets, this process comes with a trade-off. The retain accuracies also decrease, though less sharply, from around 0.55 to 0.47, indicating that the model’s resistance to MIA attacks on retained data diminishes over time. This convergence between forget and retain accuracies suggests that, by the end of the unlearning process, the model’s vulnerability to MIA attacks becomes nearly the same for both forget and retained data. While NegGrad is effective in mitigating MIA risks for forget data, it also impacts the model’s overall performance and resistance to attacks on retained data, similar to its behavior on other datasets.
Figure 16: MIA Model Accuracy on Purchase Dataset / Unlearning Algorithm NegGrad
6.3.2 SCRUB Algorithm
Figure 17 demonstrate that the SCRUB algorithm is effective in unlearning forget data for the Purchase dataset, as evidenced by the sharp decline in forget accuracy. However, this unlearning comes at a cost: both the test and validation accuracies gradually decrease, suggesting that the model’s generalization ability on unseen data is compromised as the unlearning progresses. Unlike the results seen on the CIFAR-10 and MuFac datasets, the SCRUB algorithm performs better on the Purchase dataset, showing a more successful and consistent reduction in forget accuracies while maintaining high training accuracy. Nonetheless, the trade-off in generalization performance remains a consistent challenge across datasets.
Figure 17: ML Model Accuracy on Purchase Dataset / Unlearning Algorithm SCRUB
In Figure 18, show that the SCRUB algorithm performs well in reducing MIA accuracy for forget data, making it more difficult for adversaries to identify whether forget data was part of the model’s training set. However, the model’s susceptibility to MIA attacks on retained data is also affected, albeit to a lesser extent. Overall, the SCRUB algorithm effectively balances unlearning and retaining data while mitigating MIA risks, particularly for forget data.
Figure 18: MIA Model Accuracy on Purchase Dataset / Unlearning Algorithm SCRUB
6.3.3 SFTC Algorithm
Figure 19 indicate that the SFTC algorithm is ineffective in unlearning for the Purchase dataset. While the model maintains high training accuracy, the forget accuracy remains almost unchanged, indicating that the algorithm fails to reduce the model’s ability to recall the forget data. Additionally, the test and validation accuracies remain stable, implying that the overall generalization performance is not significantly affected by the unlearning process.
Figure 19: ML Model Accuracy on Purchase Dataset / Unlearning Algorithm SFTC
In Figure 20, shows that the SFTC algorithm effectively reduces MIA attack accuracy for forget data while maintaining relatively stable performance for retained data. This suggests that the algorithm strikes a good balance between unlearning and retaining data, successfully mitigating MIA risks for forget data without significantly compromising the model’s protection of retained data. The steady separation between forget and retain accuracies highlights the algorithm’s effectiveness in reducing the model’s ability to recall forget data while preserving its resilience for retained data.
An interesting observation here is the distinction between general unlearning performance and the model’s vulnerability to MIA attacks. In Fig 19, the high forget accuracy shows that the model hasn’t fully unlearned the forget data, meaning it can still predict it accurately. However, in Fig 20, the decline in MIA attack accuracy suggests that the model becomes more resistant to adversarial attacks, likely because the features it uses to classify the forget data are becoming less distinctive. This indicates that even when the unlearning process is incomplete, the model can still improve its resilience to MIA attacks.
Figure 20: MIA Model Accuracy on Purchase Dataset / Unlearning Algorithm SFTC
6.4 Texas100 Dataset
6.4.1 NegGrad Algorithm
In Figure 21 the NegGrad algorithm achieves near complete unlearning, as evidenced by the sharp drop in forget accuracy after 20 epochs. However, this comes at the cost of the model’s overall performance, as seen by the simultaneous collapse of training, test, and validation accuracies to 0. The results suggest that while NegGrad is very effective in unlearning, it severely impacts the model’s ability to retain useful information for both training and generalization purposes. This consistent behavior across datasets, except for the Purchase, highlights a key limitation of the NegGrad algorithm: while it is highly effective in achieving unlearning, it often does so at the expense of the model’s overall performance, leading to a complete collapse in its ability to retain useful information and generalize to new data.
Figure 21: ML Model Accuracy on Texas100 Dataset / Unlearning Algorithm NegGrad
Figure 22 shows that the NegGrad algorithm is highly effective in mitigating the model’s vulnerability to MIA attacks on both forget and retained data in the Texas dataset. After 30 epochs, the model reaches a state where it no longer provides useful information to an adversary trying to conduct an MIA attack, as both the forget and retain accuracies stabilize at around 0.5, equivalent to random guessing. This suggests that while the unlearning process is complete, the model’s ability to differentiate between forget and retained data has been eliminated, achieving strong resistance to MIA attacks across the board.
Figure 22: MIA Model Accuracy on Texas100 Dataset / Unlearning Algorithm NegGrad
6.4.2 SCRUB Algorithm
[…]

### v4

@@ 7 Results
In this section, we present the results of our experiments based on the graphs we have created to demonstrate the MIA Forget and the MIA Retain accuracies throughout the unlearning epochs. We categorize the experiments in this section based on the unlearning algorithms, and for each unlearning algorithm, we present the results for the first, second, and third groups of architectures.
7.1 Negative Gradient Algorithm
7.1.1 First Group
Based on observations, (see Figure 3) the Negative Gradient algorithm demonstrated favorable performance on the CIFAR-10 dataset within the first group of target models.
(a) MIA Forget and Retain Accuracy neg grad for Cifar
(b) MIA Forget and Retain Accuracy neg grad for Mufac
(c) MIA Forget and Retain Accuracy neg grad for Purchase
(d) MIA Forget and Retain Accuracy neg grad for Texas
Figure 3: These graphs illustrate the Membership Inference Attack (MIA) performance at each epoch of the unlearning process. This specific set of results corresponds to the first group of experiments, utilizing the Negative Gradient unlearning algorithm.
Following the unlearning process (see Figure 3(a)), the MIA Retain Accuracy decreased slightly from 0.823 to 0.819, effectively maintaining the desired stability of the retain set. Simultaneously, the MIA Forget Accuracy dropped from 0.828 to 0.77. These results align with our primary objective: diminishing the MIA effectiveness on the forget set while preserving the integrity of the retain set. However, it should be noted that the MIA remained highly effective against this simple architecture, maintaining a 77% accuracy on the forget set even after unlearning. This suggests that while the Negative Gradient method provides a degree of defense against membership inference, its protective capacity in this specific scenario is relatively limited.
Regarding the unlearning process illustrated in Figure 3(b), we observe a simultaneous decline in both MIA Retain and MIA Forget accuracies. Both metrics initially plateau at 0.9, indicating a highly effective MIA attack prior to unlearning. By epoch 29, the Retain accuracy falls to 0.65 while the Forget accuracy reaches 0.478. At this stage, the model has undergone excessive unlearning; the significant degradation of the Retain accuracy suggests that the model’s predictive capacity has been partially compromised. Had the unlearning process been terminated at epoch 22, the results would have more closely aligned with our initial objectives: the Retain accuracy remained stable at 0.9, while the Forget accuracy decreased to 0.75. This indicates that while the MIA originally demonstrated high precision in detecting membership, the unlearning process successfully bolstered the protection of the target model by reducing the MIA Forget accuracy by 0.15.
Regarding the unlearning process for the Purchase-100 dataset (Figure 3(c)), both the MIA Forget and MIA Retain accuracies exhibit a downward trend. We observe a reduction of 0.075 in retain accuracy and 0.175 in forget accuracy. While this trajectory aligns with our initial hypothesis, the magnitude of these changes may not be considered a substantial enough improvement for critical real-world production environments.
Finally, regarding the Texas-100 dataset (Figure 3(d)), the MIA initially failed to achieve a high degree of effectiveness, with a baseline Retain accuracy of 0.561 and a Forget accuracy of 0.555. Following the unlearning process, we observed a slight increase in both metrics. This highlights a critical edge case: an initially ineffective MIA became paradoxically more potent due to the artifacts or ’traces’ left by the unlearning algorithm. While the increase is marginal (approximately 0.006), it demonstrates a trend where the unlearning process inadvertently enhances the model’s susceptibility to membership inference. This phenomenon suggests that unlearning can sometimes act as a signal rather than a mask, a dynamic that warrants further investigation in future work.
7.1.2 Second Group
In this subsequent series of experiments (refer to Figure 4), we replicate the epoch-wise Membership Inference Attack (MIA) evaluation. However, in this instance, the analysis is conducted on more complex model architectures to assess how increased capacity influences the unlearning process.
(a) MIA Forget and Retain Accuracy neg grad for Cifar
(b) MIA Forget and Retain Accuracy neg grad for Mufac
(c) MIA Forget and Retain Accuracy neg grad for Purchase
(d) MIA Forget and Retain Accuracy neg grad for Texas
Figure 4: These graphs illustrate the Membership Inference Attack (MIA) performance at each epoch of the unlearning process. This specific set of results corresponds to the second group of experiments, utilizing the Negative Gradient unlearning algorithm.
Regarding the CIFAR-10 dataset (see Figure 4(a)), we observe a clear declining trend in the MIA Forget Accuracy. By the 30th unlearning epoch, a simultaneous diminution of the MIA Retain Accuracy becomes apparent; however, if the unlearning process were terminated at epoch 24, we would achieve a comparatively stable MIA Retain Accuracy alongside a reduced MIA Forget Accuracy. In this instance, while the directional trend of the MIA performance aligns with our hypotheses, the absolute impact is marginal, yielding an accuracy reduction of only 0.004. Thus, while the qualitative behavior of the algorithm is correct, the quantitative impact on privacy remains limited.
Regarding the MuFac dataset (see Figure 4(b)), we observe that the MIA is largely ineffective, as the accuracy for both the forget and retain sets remains at approximately 50%. This baseline performance, equivalent to random guessing, persists throughout the unlearning process. Such results indicate that the target model does not exhibit the necessary patterns of memorization or overfitting required for a successful membership inference attack, rendering the unlearning algorithm’s impact on privacy metrics negligible in this specific scenario.
For the Purchase-100 dataset (see Figure 4(c)), the MIA accuracies for the forget and retain sets initiate at 0.634 and 0.631, respectively. Throughout the unlearning process, we observe a simultaneous diminishing trend across both datasets. This suggests that as the model’s parameters are adjusted to forget the target data, there is a collateral impact on the retain set’s membership signal, indicating that the unlearning process affects the overall posterior distribution of the model rather than isolated samples. Regarding the Texas-100 dataset (see Figure 4(d)), the results exhibit a trajectory highly similar to that of the Purchase-100 dataset. The observed trends are consistent, characterized by a significant collateral impact on the retain set. Specifically, the decline in MIA accuracy is not isolated to the forget set; rather, it extends to the retained data, suggesting that the unlearning process induces a generalized degradation of the model’s membership signal across all training samples.
7.1.3 Third Group
In this section, we present the experimental results for the third group of target models, which utilize the most complex architectures evaluated in this study. The following analysis focuses on the performance of the Negative Gradient unlearning algorithm across these high-capacity models (see Figure 5).
(a) MIA Forget and Retain Accuracy neg grad for Cifar
(b) MIA Forget and Retain Accuracy neg grad for Mufac
(c) MIA Forget and Retain Accuracy neg grad for Purchase
(d) MIA Forget and Retain Accuracy neg grad for Texas
Figure 5: These graphs illustrate the Membership Inference Attack (MIA) performance at each epoch of the unlearning process. This specific set of results corresponds to the third group of experiments, utilizing the Negative Gradient unlearning algorithm.
Regarding the CIFAR-10 dataset (Figure 5(a)), we observe a high degree of volatility in both the MIA Forget and MIA Retain accuracies during the early stages of unlearning. However, as the process progresses, a distinct downward trend emerges. Notably, the MIA Forget accuracy experiences a more pronounced decline compared to the Retain set, particularly after epoch 25. This suggests that in complex architectures, the Negative Gradient method can achieve a differentiated unlearning effect, though the overall accuracy remains high (above 0.86), indicating substantial residual memorization.
For the MUFAC dataset (Figure 5(b)), the trends for both forget and retain accuracies remain tightly coupled for the majority of the unlearning trajectory. While there is a general decrease in MIA effectiveness, the gap between the two metrics is marginal until the final epochs. By epoch 30, we see a sharp divergence where the MIA Forget accuracy drops more significantly than the Retain accuracy. This implies that for this specific dataset and architecture, extended unlearning epochs are required to achieve any meaningful separation between the membership signals of the two sets.
In the Purchase-100 experiment (Figure 5(c)), the model demonstrates high stability in its membership signal. Interestingly, the MIA Retain accuracy remains consistently higher than the Forget accuracy throughout the process. We observe a gradual, slight decline in both metrics; however, the impact is minimal (shifting less than 0.01 in accuracy). This suggests that the complex architecture utilized here is highly robust against the Negative Gradient algorithm, maintaining a nearly constant privacy profile for both sets despite the unlearning iterations.
The results for the Texas-100 dataset (Figure 5(d)) highlight a significant divergence in the latter half of the experiment. Initially, the Retain and Forget accuracies decline in tandem. However, after epoch 15, the MIA Retain accuracy begins to recover and stabilize, while the MIA Forget accuracy continues to drop sharply toward epoch 30. This represents an ideal unlearning outcome for complex models: the model successfully reduces the privacy risk associated with the forgotten data while effectively “protecting” or restoring the membership signal stability of the retained data.
In these complex architectures, the Negative Gradient algorithm shows a more nuanced behavior than in simpler models. While Texas-100 shows the most “surgical” unlearning (forgetting without harming the retain signal), Purchase-100 appears almost entirely resistant to the process. This confirms that model complexity and dataset characteristics significantly dictate the efficiency of gradient-based unlearning
7.2 Scrub Algorithm
7.2.1 First Group - Simple Target
(a) MIA Forget and Retain Accuracy scrub for Cifar
(b) MIA Forget and Retain Accuracy scrub for Mufac
(c) MIA Forget and Retain Accuracy scrub for Purchase
(d) MIA Forget and Retain Accuracy scrub for Texas
Figure 6: These graphs illustrate the Membership Inference Attack (MIA) performance at each epoch of the unlearning process. This specific set of results corresponds to the first group of experiments, utilizing the scrub unlearning algorithm.
For the CIFAR-10 dataset (Figure 6(a)), the Scrub algorithm exhibits a highly stable MIA accuracy for both sets during the initial 20 epochs, maintaining a baseline of approximately 0.83. However, after epoch 20, we observe a sharp and significant decline in the MIA Forget accuracy, which drops precipitously to nearly 0.55. While the MIA Retain accuracy also experiences a reduction, it stabilizes around 0.75. This indicates that Scrub is particularly effective at “breaking” the membership signal for the forget set in later stages, though the widening gap suggests a clear divergence between the privacy protection of the two sets.
The MUFAC results (Figure 6(b)) reveal a highly volatile unlearning trajectory. Both metrics remain at a maximum ceiling of 0.90 for the first half of the process, indicating severe vulnerability to MIA. Post-epoch 13, the MIA Forget accuracy becomes extremely unstable, fluctuating sharply before settling near 0.65. Interestingly, the Retain accuracy also undergoes significant oscillations but recovers to 0.85. This “spiky” behavior suggests that the Scrub algorithm may induce temporary structural instabilities in simpler models before reaching a state of partial unlearning.
In the Purchase-100 dataset (Figure 6(c)), we observe a distinct “reboun” effect. Immediately following the start of unlearning, both accuracies drop to a baseline of 0.50 (random chance). However, as the epochs progress, both accuracies begin to climb steadily. By epoch 30, the Retain accuracy reaches approximately 0.58, while the Forget accuracy stabilizes around 0.54. This suggests that while Scrub initially successfully masks membership, the continued unlearning process inadvertently allows the model to re-learn or expose features that slightly increase its susceptibility to MIA over time.
The results for Texas-100 (Figure 6(d)) highlight a failure to achieve the primary unlearning objective. Throughout the 30 epochs, the MIA Forget accuracy remains consistently lower than the Retain accuracy, but both show a gradual upward trend. The Retain accuracy increases from 0.56 to roughly 0.57, while the Forget accuracy oscillates around 0.555. This upward trajectory in both metrics suggests that the Scrub algorithm, when applied to this specific architecture and dataset, may actually be increasing the model’s transparency to an attacker rather than providing protection.
Compared to the Negative Gradient method, the Scrub algorithm appears more aggressive in its later stages (especially in CIFAR-10) but exhibits significant instability in MUFAC. The most concerning trend is observed in Purchase-100 and Texas-100, where unlearning leads to an increase in MIA accuracy over time. This highlights that for simple architectures, Scrub can inadvertently “leak” information as it attempts to remove it, a phenomenon that emphasizes the importance of early stopping at the “dip” (e.g., epoch 2 in Purchase-100).
7.2.2 Second Group
(a) MIA Forget and Retain Accuracy scrub for Cifar
(b) MIA Forget and Retain Accuracy scrub for Mufac
(c) MIA Forget and Retain Accuracy scrub for Purchase
(d) MIA Forget and Retain Accuracy scrub for Texas
Figure 7: These graphs illustrate the Membership Inference Attack (MIA) performance at each epoch of the unlearning process. This specific set of results corresponds to the second group of experiments, utilizing the scrub unlearning algorithm.
In the CIFAR-10 experiment (Figure 7(a)), the Scrub algorithm demonstrates a notable “plunge” in MIA accuracy during the initial five epochs. Both the Forget and Retain accuracies drop sharply from a baseline of 0.75, with the Forget accuracy reaching a local minimum near 0.60. Following this initial phase, the Retain accuracy recovers and stabilizes around 0.675, while the Forget accuracy continues a steady, gradual decline toward 0.575. This represents an effective unlearning trajectory, as the algorithm successfully creates a consistent gap between the membership signals of the two sets after the initial instability.
The results for the MUFAC dataset (Figure 7(b)) highlight a scenario where the MIA is initially ineffective, with both accuracies hovering at a baseline of 0.50 (random chance). However, as the unlearning process continues, the model exhibits significant instability. We observe frequent “spikes” and fluctuations in both metrics, particularly between epochs 15 and 30. Despite these oscillations, the accuracies generally return to the 0.50 baseline. This suggests that while Scrub does not increase overall privacy risk here, it introduces transient structural changes that temporarily make certain samples more identifiable before they are re-masked.
For the Purchase-100 dataset (Figure 7(c)), the Scrub algorithm shows a distinct “delayed reaction.” For the first 12 epochs, both MIA Forget and Retain accuracies remain high and stable at approximately 0.635. At epoch 13, there is a sudden and significant drop in both metrics. While the Retain accuracy partially recovers to approximately 0.595, the Forget accuracy continues to decline, ending near 0.565. This indicates that the unlearning mechanism for this architecture requires a specific threshold of iterations before the membership signal is successfully disrupted.
The Texas-100 results (Figure 7(d)) demonstrate a more linear and “surgical” unlearning process compared to the other datasets in this group. We observe a consistent downward trend for both metrics from the outset. Crucially, the MIA Forget accuracy declines at a steeper rate than the Retain accuracy throughout the 30 epochs. By the end of the process, the Forget accuracy reaches a low of 0.57, while the Retain accuracy remains higher at roughly 0.62. This suggests that for intermediate architectures on the Texas dataset, Scrub effectively prioritizes the removal of the forget set’s membership information while preserving a stronger signal for the retained data.
For this second group of models, the Scrub algorithm appears more stable than it was with simpler architectures, particularly in the Texas-100 and CIFAR-10 cases. The most prominent feature of this group is the initial volatility (CIFAR-10) and delayed disruption (Purchase-100), suggesting that the timing of the “unlearning shock” is highly dependent on the dataset’s characteristics. Unlike the first group, we do not see a significant “rebound” where accuracies increase at the end, implying that higher model complexity may help in maintaining the privacy gains achieved during the unlearning process.
7.2.3 Third Group
(a) MIA Forget and Retain Accuracy scrub for Cifar
(b) MIA Forget and Retain Accuracy scrub for Mufac
(c) MIA Forget and Retain Accuracy scrub for Purchase
(d) MIA Forget and Retain Accuracy scrub for Texas
Figure 8: These graphs illustrate the Membership Inference Attack (MIA) performance at each epoch of the unlearning process. This specific set of results corresponds to the third group of experiments, utilizing the scrub unlearning algorithm.
For the CIFAR-10 (Figure 8(a)) dataset using complex architectures, the Scrub algorithm demonstrates a highly effective and stable unlearning profile. While the MIA Retain Accuracy remains remarkably consistent, fluctuating minimally around 0.88, the MIA Forget Accuracy exhibits a steady and significant decline from 0.875 down to approximately 0.835 by epoch 30. This clear divergence indicates that for high-capacity models, Scrub can successfully target and diminish the membership signal of the forget set without compromising the privacy profile of the retained data.
In the MUFAC experiment (Figure 8(b)), we observe a similar but more pronounced downward trajectory for the forget set. The MIA Forget Accuracy starts near 0.86 and decreases consistently, reaching a low of approximately 0.79 by the end of the unlearning process. Meanwhile, the MIA Retain Accuracy maintains a high baseline around 0.86, showing only minor oscillations. This represents a successful unlearning scenario where the “forgetting” is surgical, specifically reducing the identifiable traces of the target samples in a high-dimensional model space.
The results for the Purchase-100 dataset (Figure 8(c)) reveal a high degree of volatility but a relatively stable overall privacy baseline. Both the forget and retain accuracies oscillate significantly between 0.790 and 0.806. Despite these fluctuations, there is no significant downward trend for either metric, with the accuracies ending near their initial starting points. This suggests that for complex architectures on Purchase-100, the Scrub algorithm struggles to find a consistent gradient to effectively reduce the membership signal, resulting in a model that remains equally susceptible to MIA throughout the process.
The Texas-100 dataset (Figure 8(d)) presents a paradoxical result. Following the start of unlearning, both MIA Forget and Retain accuracies actually increase during the first 15 epochs, rising from 0.786 to nearly 0.792. In the latter half of the process, both metrics begin to decline, eventually returning to near-baseline levels. This “bell-curve behavior suggests that for complex Texas-100 models, the Scrub algorithm initially makes the data points more identifiable—likely due to model updates creating distinctive artifacts—before finally beginning to mask the membership signal in the final epochs.
In this final group, the Scrub algorithm shows its greatest strength in CIFAR-10 and MUFAC, where it achieves a clear separation between forget and retain membership signals. However, the Texas-100 results serve as a warning: in complex architectures, the unlearning process can initially increase privacy risk before reducing it. Furthermore, the resistance observed in Purchase-100 indicates that even advanced algorithms like Scrub are not universally effective and are highly dependent on the underlying data distribution and model capacity.
7.3 SFTC Algorithm
7.3.1 First Group
(a) MIA Forget and Retain Accuracy SFTC for Cifar
(b) MIA Forget and Retain Accuracy SFTC for Mufac
(c) MIA Forget and Retain Accuracy SFTC for Purchase
(d) MIA Forget and Retain Accuracy SFTC for Texas
Figure 9: These graphs illustrate the Membership Inference Attack (MIA) performance at each epoch of the unlearning process. This specific set of results corresponds to the first group of experiments, utilizing the sftc unlearning algorithm.
For the CIFAR-10 dataset (Figure 9(a)), the SFTC algorithm demonstrates high precision in selectively targeting the forget set. Initially, the MIA Forget Accuracy starts near 0.62, but it experiences a sharp and consistent decline, reaching a low of approximately 0.46 by epoch 30—falling below the threshold of random guessing. Crucially, the MIA Retain Accuracy remains exceptionally stable throughout the process, maintaining a baseline above 0.80. This suggests that SFTC effectively removes membership traces of the forget set while preserving the model’s structural knowledge of the retained data.
The MUFAC results (Figure 9(b)) highlight a highly volatile unlearning process characterized by significant oscillations. While both metrics start at a high baseline of 0.85, the MIA Forget Accuracy undergoes dramatic “spikes,” momentarily dropping to 0.70 before rebounding. By the final epochs, it stabilizes significantly lower than its starting point, around 0.70. The MIA Retain Accuracy also fluctuates but stays generally higher, around 0.85. This indicates that while SFTC manages to reduce privacy risk, the unlearning path for MUFAC in simple models is unstable, likely due to the sensitive nature of the dataset’s features.
In the Purchase-100 experiment (Figure 9(c)), we observe a steady downward trend for both the forget and retain sets. The MIA Forget Accuracy starts at 0.635 and concludes at approximately 0.605. However, the MIA Retain Accuracy also mirrors this decline, falling from 0.635 to roughly 0.625. While a gap between the two metrics eventually emerges, the simultaneous decline suggests that the SFTC algorithm has a “thinning” effect on the model’s overall confidence, affecting both sets rather than performing a purely surgical removal of the forget set’s identity.
Regarding the Texas-100 dataset (Figure 9(d)), the SFTC algorithm appears to struggle with providing a significant defense. Both the forget and retain accuracies remain highly coupled and volatile throughout the 30 epochs, hovering between 0.557 and 0.573. There is no distinct downward trajectory for the forget set; in fact, both accuracies end near their starting points. This lack of impact suggests that for simple architectures on Texas-100, the teacher-student dynamics of SFTC fail to effectively distinguish or “scrub” the membership signals, resulting in negligible privacy gains.
Comparing the SFTC algorithm to the previous Scrub and Negative Gradient methods on simple architectures reveals a key strength: selectivity. In datasets like CIFAR-10, SFTC achieves the most surgical separation between forget and retain signals seen thus far. However, its performance on Texas-100 remains a point of concern, showing that gradient-based and teacher-student unlearning both face difficulties when the initial membership signal is weak or the data distribution is high-dimensional.
7.3.2 Second Group
(a) MIA Forget and Retain Accuracy SFTC for Cifar
(b) MIA Forget and Retain Accuracy SFTC for Mufac
(c) MIA Forget and Retain Accuracy SFTC for Purchase
(d) MIA Forget and Retain Accuracy SFTC for Texas
Figure 10: These graphs illustrate the Membership Inference Attack (MIA) performance at each epoch of the unlearning process. This specific set of results corresponds to the second group of experiments, utilizing the sftc unlearning algorithm.
In the CIFAR-10 scenario (Figure 10(a)), the SFTC algorithm exhibits a highly effective unlearning trajectory. The MIA Forget Accuracy starts at approximately 0.55 and undergoes a steady, consistent decline, reaching a final accuracy of 0.42 by epoch 30. Simultaneously, the MIA Retain Accuracy shows a significant upward trend, rising from 0.57 to approximately 0.80. This suggests that as the model selectively forgets the target data, its generalization and membership signal for the retained data are strengthened. This raise the privacy for the retain data.
For the MUFAC dataset(Figure 10(b)), the MIA attack remains largely ineffective, with accuracies hovering near the random guessing baseline of 0.50. We observe minor fluctuations where both the forget and retain accuracies rise slightly to 0.502 before stabilizing back at 0.50. This indicates that for this intermediate architecture, the membership signals are naturally well-masked, and the SFTC unlearning process maintains this high level of privacy without introducing new identifiable artifacts.
[…]